# Computer-Assisted Proofs of Gap Solitons in Bose-Einstein Condensates

### v2

@@ 2.1 Solving for The First Order Coefficients: The Stable Bundle
Although $\lambda$ is expected to be real, we perform computations in the complex space $X_{\scriptscriptstyle{\mathsf{F}}}$ because the sequence space components involve complex coefficients, and our numerical methods use complex floating-point vectors. Once the existence of a solution in the complex space is established, we demonstrate that $\lambda$ is real and satisfies $\lambda<0$.
Observe that if $(\lambda,v)$ solves (23), then any scalar multiple of $v$ is also a solution. To guarantee uniqueness, which is essential for the Newton-Kantorovich approach used in Theorem 2, we introduce a phase condition as an additional equation.
@@
A convenient formulation of the phase condition involves defining the sequence $\mathbf{1}_{M}$, indexed over $\mathbb{Z}$, such that $(\mathbf{1}_{M})_{m}=1$ for $|m|\leq M$ and $(\mathbf{1}_{M})_{m}=0$ for $|m|>M$. This allows the phase condition to be expressed as the dot product $\eta(v)=\mathbf{1}_{M}\cdot v^{(1)}$. With this normalization condition on $v$, we define the validation map $F$ by
@@
We adopt the following conventions regarding the action of truncation operator on elements $(\lambda,v)$ of the product space $X_{\scriptscriptstyle{\mathsf{F}}}$
@@
$[P,Q]\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\{m\in\mathbb{Z}:P\leq|m|\leq Q\right\},\quad(P,Q]\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\{m\in\mathbb{Z}:P<|m|\leq Q\right\},\$
@@
We define the support of a sequence $p$ as the set $R$ such that $p_{m}=0$ for all $m\notin R$. Sequences with finite support can be represented as finite-dimensional vectors, which are suitable for computational manipulation. We refer to such sequences as computable sequences. In contrast, sequences whose support lies within the interval $(M,\infty)$ for some $M\in\mathbb{N}$ are infinite-dimensional and cannot naturally be represented as finite-dimensional vectors. These sequences are referred to as infinite tails. With this terminology established, we proceed to describe the subsequent steps of our method, introducing the element $\bar{x}$ and the operator $A$ required to apply Theorem 2. Let $\bar{x}=\bar{x}_{\scriptscriptstyle{\mathsf{F}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(\bar{\lambda},\bar{v})$ denote an element of $X_{{\scriptscriptstyle{\mathsf{F}}}}$ with finite support, which is
@@
For all $m\in[0,2M]$, $r_{m}=0$. Now, for $m>2M$, we have $k-m\leq-M$. If $m<-2M$, then $k-m>3M$. Therefore for $m\in(2M,\infty)$ we have $p_{k-m}=0$. The argument is analogous when $-M<k\leq 0$. Therefore we have that $\left(p*_{\scriptscriptstyle{\mathsf{F}}}r\right)_{k}=0$ for all $k\in[0,M]$. ∎
@@
We now proceed to derive the bound $Z_{1}=Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})$ that satisfies (8).
@@
$\|[DF(\bar{x}_{\scriptscriptstyle{\mathsf{F}}}+\delta)-DF(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})]h\|_{X_{\scriptscriptstyle{\mathsf{F}}}}=\|(0,\lambda v_{\delta}-\lambda_{\delta}v)\|_{X_{\scriptscriptstyle{\mathsf{F}}}}\leq 2r$
@@
$r_{\scriptscriptstyle{\mathsf{F}}}=\frac{1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})-\sqrt{(1-Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}-2Y(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})}}{Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})}.$
@@
where given a complex number $z\in\mathbb{C}$, $z^{*}$ denotes its complex conjugate. For any $x\in X_{\scriptscriptstyle{\mathsf{F}}}$, we have
@@ 2.2 Solving for Higher Order Coefficients
$[B^{(i)}(w,v)]_{n,m}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\begin{cases}w_{0,m}^{(i)}-\gamma^{(i)}_{m},&n=0,m\in\mathbb{Z}\\ w_{1,m}^{(i)}-v^{(i)}_{m},&n=1,m\in\mathbb{Z}\\ 0,&n\geq 2,m\in\mathbb{Z},\end{cases}$ (45)
@@
$[B^{(i)}(\bar{w},v)]_{n,m}=\begin{cases}\bar{v}_{m}^{(i)}-v^{(i)}_{m},&n=1,m\in\mathbb{Z}\\ 0,&n\neq 1,m\in\mathbb{Z}\end{cases}\quad\text{for}\quad i=1,2.$ (49)
@@
$\prod\limits_{\scriptscriptstyle{[0,N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(N,3N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(3N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}-A_{\infty}DF(\bar{w})\left[\prod\limits_{\scriptscriptstyle{[0,3N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(3N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}\right]$
@@
$[Df(\bar{w})-Df(\bar{w}+\mathbf{\delta})]h=\begin{pmatrix}0\\ -3c\left(2\bar{w}^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}h^{(1)}+\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}h^{(1)}\right)\\ \end{pmatrix},$
@@
$\lVert Az\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}$ $\leq\lVert A_{f}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}})}+\lVert\prod\limits_{\scriptscriptstyle{[2,N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}L_{\lambda}^{-1}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}+\lVert\prod\limits_{\scriptscriptstyle{(N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}L_{\lambda}^{-1}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}.$
@@
Fix parameters $a$, $b$, and $c$, a weight $\nu$ for the norm in (15). We consider a solution of the bundle sequence equation (23) of the form $x_{\scriptscriptstyle{\mathsf{F}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(\lambda,v^{(1)},v^{(2)},0,0)$ in the space ×$\mathbb{C}\times S^{4}_{\scriptscriptstyle{\mathsf{F}}}$ where $(\lambda,v^{(1)},v^{(2)})$ is in the ball $B_{X_{\scriptscriptstyle{\mathsf{F}}}}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}},r_{\scriptscriptstyle{\mathsf{F}}})$. Suppose we have $\bar{w}$ in $X_{\scriptscriptstyle{\mathsf{TF}}}$ such that
@@
$Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<1\quad\text{and}\quad Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})<min\left(\frac{(1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}}{2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})},r_{\scriptscriptstyle{\mathsf{TF}}}^{*}\right)$
@@
$r_{\scriptscriptstyle{\mathsf{TF}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\frac{1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})-\sqrt{(1-Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*}))^{2}-2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})}}{Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})}$
@@ 3 Solving the Boundary-Value Problem
$u_{2}(-1)=0,\quad u_{3}(-1)=1,\quad u_{4}(-1)=0,\quad u(1)=W_{1}(\theta,\sigma),\quad u(1)=W_{2}(\theta,\sigma).$
We have 5 boundary conditions for the 4 components of the solution. To balance the system we will solve for $\sigma$ and fix the value of $\theta$ and $\mathsf{L}$ We now consider the following Chebyshev expansion for the solution of the BVP (57)
$s^{(i)}(t)=s^{(i)}_{0}+2\sum_{m\geq 1}s^{(i)}_{m}T_{m}(t),\quad i=1,2,3,4,$
where $T_{m}:[-1,1]\to\mathbb{R}$ with $m\geq 0$ are the Chebyshev polynomials of the first kind. Given a weight $\omega\geq 1$ we define the sequence space $S_{\scriptscriptstyle{\mathsf{C}}}$ of Chebyshev coefficients as
@@
For any two sequences $u,v\in S_{\scriptscriptstyle{\mathsf{C}}}$, we define their discrete convolution $*_{\scriptscriptstyle{\mathsf{C}}}:S_{\scriptscriptstyle{\mathsf{C}}}\times S_{\scriptscriptstyle{\mathsf{C}}}\to S_{\scriptscriptstyle{\mathsf{C}}}$ by
@@
Since we are working with boundary conditions, for any sequence $s\in S_{\scriptscriptstyle{\mathsf{C}}}$ we will use the following notation when we refer to the sequence as a function
@@
In particular the evaluations at $-1$ and $1$ are given by
@@
We proceed with the definition of the validation map for the soliton boundary-value problem. Notably, we employ the same notation as in previous contexts to highlight the structural parallels in our approach. As derived in Section 2, let $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ be a parameterization of the local stable manifold $W^{s}_{\rm loc}(\gamma)$ associated with the periodic orbit $\gamma$. The validation map is defined as follows:
@@
$\prod\limits_{\scriptscriptstyle{[0,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}B(x)=\prod\limits_{\scriptscriptstyle{\{0\}}}^{\mathsf{\scriptscriptstyle{C}}}B(x)$ =$\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left(s^{(1)}(1)-W^{(1)}(\theta,\sigma),s^{(2)}(1)-W^{(2)}(\theta,\sigma),s^{(3)}(-1)-1,s^{(4)}(-1)\right),$
$\prod\limits^{\mathsf{\scriptscriptstyle{R}}}B(x)$ =$\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(s^{(2)}(-1),0).$
We now consider $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}=(\bar{\sigma},\bar{s})$ an element in $X_{\scriptscriptstyle{\mathsf{C}}}$ supported in $[0,M)$ and such that $-1<\bar{\sigma}<1$. We define an approximate inverse derivative $A:X_{\scriptscriptstyle{\mathsf{C}}}\to X_{\scriptscriptstyle{\mathsf{C}}}$ as follows
@@
where $A_{f}$ can be represented as a $[4(M+1)+1]\times[4(M+1)+1]$ matrix that in practice corresponds to a numerical inverse of the truncated derivative of $F$ evaluated at $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}$. We suppose that for a fixed set of parameters $a$, $b$ and $c$ there exists $\bar{x}_{\scriptscriptstyle{\mathsf{TF}}}=(\bar{w}^{(1)},\bar{w}^{(2)},\bar{w}^{(3)},\bar{w}^{(4)})\in S^{4}_{\scriptscriptstyle{\mathsf{TF}}}$ with finite support such that there exists a parameterization of the stable manifold $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ attached to the periodic orbit $\gamma$ satisfying the inequality in equation (55). This is the parameterization given by Theorem 11. We are now ready to provide computable bounds as those required by Theorem 2.

### v3

@@ 2.1 Solving for The First Order Coefficients: The Stable Bundle
Although $\lambda$ is expected to be real, we perform computations in the complex space $X_{\scriptscriptstyle{\mathsf{F}}}$ because the sequence space components involve complex coefficients, and our numerical methods use complex floating-point vectors. Once the existence of a solution in the complex space is established, we demonstrate that $\lambda$ is real and satisfies $\lambda<0$. Observe that if $(\lambda,v)$ solves (25), then any scalar multiple of $v$ is also a solution. To guarantee uniqueness, which is essential for the Newton-Kantorovich approach used in Theorem 2, we introduce a phase condition as an additional equation:
@@
The phase condition sets the first component of $v(0)$ equal to $l$. A convenient formulation of the phase condition involves defining the sequence $\mathbf{1}_{M}$, indexed over $\mathbb{Z}$, such that $(\mathbf{1}_{M})_{m}=1$ for $|m|\leq M$ and $(\mathbf{1}_{M})_{m}=0$ for $|m|>M$. This allows the phase condition to be expressed as the dot product $\eta(v)=\mathbf{1}_{M}\cdot v^{(1)}$. With this normalization condition on $v$, we define the validation map $F$ by
@@
We adopt the following conventions regarding the action of truncation operator for a set of indices $R$ on elements $(\lambda,v)$ of the product space $X_{\scriptscriptstyle{\mathsf{F}}}$
@@
$P,Q]\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\{m\in\mathbb{Z}:P\leq|m|\leq Q\right\},\quad(P,Q]\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left\{m\in\mathbb{Z}:P<|m|\leq Q\right\},\\$
@@
We define the support of a sequence $p$ as the set of indices $R$ such that $p_{m}\neq 0$ for all $m\in R$. Sequences with finite support can be represented as finite-dimensional vectors, which are suitable for computational manipulation. We refer to such sequences as computable sequences. In contrast, sequences whose support lies within the interval $(M,\infty)$ for some $M\in\mathbb{N}$ are infinite-dimensional and cannot naturally be represented as finite-dimensional vectors. These sequences are referred to as infinite tails. With this terminology established, we proceed to describe the subsequent steps of our method, introducing the element $\bar{x}$ and the operator $A$ required to apply Theorem 2. Let $\bar{x}=\bar{x}_{\scriptscriptstyle{\mathsf{F}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(\bar{\lambda},\bar{v})$ denote an element of $X_{{\scriptscriptstyle{\mathsf{F}}}}$ with finite support, which is
@@
Even if our notation based on truncation operators may appear cumbersome at first, our choice is deliberate. It is designed so that the bounds and operations in the paper closely match those implemented in the computer code, making the translation from paper to code as transparent as possible and simplifying the verification of our implementation.
@@
For all $m\in[0,2M]$, $r_{m}=0$. Now, for $m>2M$, we have $k-m\leq-M$. If m<2Mm<-2M, then km>3Mk-m>3M. Therefore for $m\in(2M,\infty)$ we have $p_{k-m}=0$. The argument is analogous when $M<k\leq 0$. Therefore we have that $\left(p*_{\scriptscriptstyle{\mathsf{F}}}r\right)_{k}=0$ for all $k\in[0,M]$. ∎
@@
We now proceed to derive the bound $Z_{1}=Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})$ that satisfies (8). In what follows, we write $B(X_{\scriptscriptstyle{\mathsf{F}}})$ for the space of continuous linear operators on $X_{\scriptscriptstyle{\mathsf{F}}}$.
@@
$\|[DF(\bar{x}_{\scriptscriptstyle{\mathsf{F}}}+\delta)-DF(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})]h\|_{X_{\scriptscriptstyle{\mathsf{F}}}}=\|(0,\lambda v_{\delta}+\lambda_{\delta}v)\|_{X_{\scriptscriptstyle{\mathsf{F}}}}\leq 2r$
@@
$r_{\scriptscriptstyle{\mathsf{F}}}=\frac{1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})-\sqrt{(1-Z_{1}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}-2Y(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})}}{Z_{2}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}})}.$
@@
where given a complex number $\in\mathbb{C}, zz^{*}$ denotes its complex conjugate. For any $x\in X_{\scriptscriptstyle{\mathsf{F}}}$, we have
@@ 2.2 Solving for Higher Order Coefficients
$B^{(i)}(w,v)]_{n,m}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\begin{cases}w_{0,m}^{(i)}-\gamma^{(i)}_{m},&n=0,m\in\mathbb{Z}\\ w_{1,m}^{(i)}-v^{(i)}_{m},&n=1,m\in\mathbb{Z}\\ 0,&n\geq 2,m\in\mathbb{Z},\end{cases}$ (47)
@@
$B^{(i)}(\bar{w},v)]_{n,m}=\begin{cases}\bar{v}_{m}^{(i)}-v^{(i)}_{m},&n=1,m\in\mathbb{Z}\\ 0,&n\neq 1,m\in\mathbb{Z}\end{cases}\quad\text{for}\quad i=1,2.$ (51)
@@
$\hskip-28.45274pt\prod\limits_{\scriptscriptstyle{[0,N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(N,3N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(3N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}-A_{\infty}DF(\bar{w})\left[\prod\limits_{\scriptscriptstyle{[0,3N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(3M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}+\prod\limits_{\scriptscriptstyle{(3N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}\right]$
@@
$Df(\bar{w})-Df(\bar{w}+\mathbf{\delta})]h=\begin{pmatrix}0\\ -3c\left(2\bar{w}^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}h^{(1)}+\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}\delta^{(1)}*_{\scriptscriptstyle{\mathsf{TF}}}h^{(1)}\right)\\ \end{pmatrix}$,
@@
$\lVert Az\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}$ $\leq\lVert A_{f}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}+\lVert\prod\limits_{\scriptscriptstyle{[2,N]}}^{\mathsf{\scriptscriptstyle{T}}}\prod\limits_{\scriptscriptstyle{(M,\infty)}}^{\mathsf{\scriptscriptstyle{F}}}L_{\lambda}^{-1}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}+\lVert\prod\limits_{\scriptscriptstyle{(N,\infty)}}^{\mathsf{\scriptscriptstyle{T}}}L_{\lambda}^{-1}z\rVert_{X_{\scriptscriptstyle{\mathsf{TF}}}}.$
@@
Fix parameters $a$, $b$, and $c$, a weight $\nu$ for the norm in (17). We consider a solution of the bundle sequence equation (25) of the form $x_{\scriptscriptstyle{\mathsf{F}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(\lambda,v^{(1)},v^{(2)},0,0)$ in the space $\mathbb{C}\times S^{4}_{\scriptscriptstyle{\mathsf{F}}}$ where $(\lambda,v^{(1)},v^{(2)})$ is in the ball $B_{X_{\scriptscriptstyle{\mathsf{F}}}}(\bar{x}_{\scriptscriptstyle{\mathsf{F}}},r_{\scriptscriptstyle{\mathsf{F}}})$. Suppose we have $\bar{w}$ in $X_{\scriptscriptstyle{\mathsf{TF}}}$ such that
@@
$Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})<1\quad\text{and}\quad Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})<\frac{(1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}}))^{2}}{2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})}$
@@
$r_{\scriptscriptstyle{\mathsf{TF}}}\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\frac{1-Z_{1}(\bar{w},\bar{x}_{\scriptscriptstyle{\mathsf{F}}})-\sqrt{(1-Z_{1}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*}))^{2}-2Y(\bar{w},r_{\scriptscriptstyle{\mathsf{F}}})Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})}}{Z_{2}(\bar{w},r_{\scriptscriptstyle{\mathsf{TF}}}^{*})}$
@@ 3 Solving the Boundary-Value Problem
$u_{2}(-1)=0,\quad u_{3}(-1)=1,\quad u_{4}(-1)=0,\quad u_{1}(1)=W_{1}(\theta,\sigma),\quad u_{2}(1)=W_{2}(\theta,\sigma).$
We have 5 boundary conditions for the 4 components of the solution. To balance the system we will solve for $\sigma$ and fix the value of $\theta$ and $\mathsf{L}$. We now consider the following Chebyshev expansion for the solution of the BVP (59)
$s^{(i)}(t)=s^{(i)}_{0}+2\sum_{m\geq 1}s^{(i)}_{m}T_{m}(t),\quad i=1,2,3,4,$ (60)
where $T_{m}:[-1,1]\to\mathbb{R}$ with $m\geq 0$ are the Chebyshev polynomials of the first kind. These polynomials are defined recursively as follows:
$T_{0}(t)=1,\quad T_{1}(t)=t,\quad T_{m+1}(t)=2tT_{m}(t)-T_{m-1}(t)\quad\text{for }m\geq 1.$
Given a weight $\omega\geq 1$ we define the sequence space $S_{\scriptscriptstyle{\mathsf{C}}}$ of Chebyshev coefficients as
@@
For any two sequences $u,v\in S_{\scriptscriptstyle{\mathsf{C}}}$, we define their discrete convolution $_{\scriptscriptstyle{\mathsf{C}}}:S_{\scriptscriptstyle{\mathsf{C}}}\times S_{\scriptscriptstyle{\mathsf{C}}}\to S_{\scriptscriptstyle{\mathsf{C}}}$ by
@@
For any sequence $s\in S_{\scriptscriptstyle{\mathsf{C}}}$ we will use the following notation when we refer to the sequence as a function
@@
Since Chebyshev polynomials satisfy the identity $T_{k}(\cos\theta)=\cos(k\theta)$, we have that $T_{k}(-1)=(-1)^{k}$ and $T_{k}(1)=1$ for all $k$. Therefore, the evaluations at 1-1 and $1$ of a sequence are given by
@@
We proceed with the definition of the validation map for the soliton boundary-value problem. We employ the same notation as in previous sections to highlight the structural parallels in our approach. Moreover, as derived in Section 2, let $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ be a parameterization of the local stable manifold $W^{s}_{\rm loc}(\gamma)$ associated with the periodic orbit $\gamma$.
In this case, the sequence equation follows the derivation in [21]. For completeness, we briefly outline the argument for the first component of $s$, which satisfies $s^{(1)}(1)=W^{(1)}(\theta,\sigma)$. We begin by integrating the differential equation $\dot{s}(t)=f(s(t))$ from $t$ to $1$:
$s^{(1)}(1)-s^{(1)}(t)=\kappa\int_{t}^{1}f^{(1)}(s(u))\,du.$ (64)
We then substitute in (64) the Chebyshev expansion of the solution, given in equation (60), along with the Chebyshev expansion of the vector field $f$:
$f^{(i)}(s(t))=\phi^{(i)}_{0}+2\sum_{m\geq 1}\phi^{(i)}_{m}T_{m}(t),\quad i=1,2,3,4.$
To evaluate the integral in (64), we use the recurrence relation for integrals of Chebyshev polynomials:
$\int T_{m}(t)\,dt=\frac{1}{2}\left(\frac{T_{m+1}(t)}{m+1}-\frac{T_{m-1}(t)}{m-1}\right),\quad\text{for }m\geq 2.$
Matching Chebyshev coefficients on both sides yields a system of equations for the unknown coefficients $a_{m}$. This leads to the following equations for the Chebyshev coefficients:
$s_{0}+2\sum_{m\geq 1}(-1)^{m}s_{m}-W^{(1)}(\theta,\sigma)=0,$
$2m\,s_{m}^{(1)}+\phi^{(1)}_{m+1}-\phi^{(1)}_{m-1}=0,\quad\text{for }m\geq 1.$
Since the Chebyshev expansion of a sum or difference of two functions is obtained by adding or subtracting their Chebyshev coefficients, and the Chebyshev coefficients of a product are given by the discrete Chebyshev convolution (63), any polynomial in the components of $s$ can be written using sums and discrete convolutions of their Chebyshev coefficients. In particular, the Chebyshev coefficients of $f(s)$ can be expressed directly in terms of those of $s$.
Applying the same procedure to the remaining three components of $s$, we obtain a complete sequence equation. Further details on the derivation of the sequence equations can be found in [21]. This construction defines our validation map.
@@
$\prod\limits_{\scriptscriptstyle{[0,\infty)}}^{\mathsf{\scriptscriptstyle{C}}}B(x)=\prod\limits_{\scriptscriptstyle{\{0\}}}^{\mathsf{\scriptscriptstyle{C}}}B(x)$ $\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,\left(s^{(1)}(1)-W^{(1)}(\theta,\sigma),s^{(2)}(1)-W^{(2)}(\theta,\sigma),s^{(3)}(-1)-1,s^{(4)}(-1)\right),$
$\prod\limits^{\mathsf{\scriptscriptstyle{R}}}B(x)$ $\,\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\,(s^{(2)}(-1),0).$
We now consider $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}=(\bar{\sigma},\bar{s})$ an element in $X_{\scriptscriptstyle{\mathsf{C}}}$ supported in $[0,M)$ and such that $1<\bar{\sigma}<1$. We define an approximate inverse derivative $A:X_{\scriptscriptstyle{\mathsf{C}}}\to X_{\scriptscriptstyle{\mathsf{C}}}$ as follows
@@
where $A_{f}$ can be represented as a $4(M+1)+1]\times[4(M+1)+1]$ matrix that in practice corresponds to a numerical inverse of the truncated derivative of $F$ evaluated at $\bar{x}_{\scriptscriptstyle{\mathsf{C}}}$. We suppose that for a fixed set of parameters $a$, $b$ and $c$ there exists $\bar{x}_{\scriptscriptstyle{\mathsf{TF}}}=(\bar{w}^{(1)},\bar{w}^{(2)},\bar{w}^{(3)},\bar{w}^{(4)})\in S^{4}_{\scriptscriptstyle{\mathsf{TF}}}$ with finite support such that there exists a parameterization of the stable manifold $W:[-1,1]\times[0,2\pi]\to\mathbb{R}^{4}$ attached to the periodic orbit $\gamma$ satisfying the inequality in equation (57). This is the parameterization given by Theorem 11. We are now ready to provide computable bounds as those required by Theorem 2.